# The Generation Phases of Flow Matching: a Denoising Perspective

### v1

@@ Appendix A Details on the losses
where $\sigma_{\max}$ is typically around $0.5$. From Remark 1, classical and generative denoisers are equivalent up to the reparameterization
$t(\sigma)=\frac{1}{1+\sigma}.$
A change of variables in (13) gives:
@@
Appendix B Details on training generation and metrics
All networks are trained with the same random initialization to ensure comparability. For CIFAR-10 we train for 400 epochs with batch size $128$, and for CelebA-64 we train for 300 epochs with batch size $128$. We apply exponential moving average to stabilize training. For CelebA-64, we use a U-Net architecture [Ronneberger et al., 2015] as in Ho et al. [2020]. For CIFAR-10, we adopt the architecture from the torchcfm library [Tong et al., 2024].
The 10 denoisers of Figure 2 are trained independently with 400 epochs each. All samples are generated by solving the ODE using the dopri5 scheme from $t=0$ to $t=1$ with $100$ timesteps; generating samples using a denoiser $D$ means that using the velocity $v(x,t)=\tfrac{D(x,t)-x}{1-t}$.
To measure generation quality, we use the standard Fréchet Inception Distance [Heusel et al., 2017] with the Inception-V3 features [Szegedy et al., 2016]; as recommended by Stein et al. [2023] we also compute it with the DINOv2 embedding.
Appendix C Additional results on CIFAR10 and CelebA-64
We report additional results on CIFAR10 and CelebA-64. In particular, we broaden the range of weights considered by including those defined in Table 2, with weights putting more or less emphasis than $w_{t}^{\text{FM}}$ on large times ($w_{t}^{\text{pow 1}}$, $w_{t}^{\text{pow 3}}$) and one putting more emphasis on $t=0.2$, $w_{t}^{\text{mid-0.2}}$.
Tables 3 and 4 present the performance of all models in terms of PSNR and the two geenrative metrics, FID and DINO. Figure 7 provides a visual overview for CelebA-64, as was done for CIFAR-10 in the main paper.
Losses $\mathcal{L}$ Parametrization classes $\mathcal{C}$
$\mathcal{L}_{\text{classic}}:\,w_{t}^{\text{classic}}=\frac{1}{t^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$: $D(x,t)=x+(1-t)N^{\theta}(x,t)$
$\mathcal{L}_{\text{den}}:\,w_{t}^{\text{den}}=1$
$\mathcal{L}_{\text{pow 1}}:\,w_{t}^{\text{pow 1}}=\frac{1}{(1-t)}$
$\mathcal{L}_{\text{pow 3}}:\,w_{t}^{\text{pow 3}}=\frac{1}{(1-t)^{3}}$
$\mathcal{L}_{\text{mid}}:\,w_{t}^{\text{mid}}=\frac{1}{(0.5-t)^{2}}$
$\mathcal{L}_{\text{mid-0.2}}:\,w_{t}^{\text{mid-0.2}}=\frac{1}{(0.2-t)^{2}}$
Table 3: PSNR and FID for the different losses, to be compared with the standard FM (bottom line, corresponding to loss $\mathcal{L}_{\text{FM}}$ and parametrization $\mathcal{C}_{\mathrm{I+NN}}$) . PSNR computed on 1000 images; FID on 10k test images; CIFAR-10, 400 epochs.
Loss Class PSNR (↑) FID / DINO (↓)
@@ Appendix C Additional results on CIFAR10 and CelebA-64
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 14.41 18.16 23.42 32.89 37.33 12.97 / 506.54
$\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\text{NN}}$ 14.42 18.17 23.35 32.54 36.70 26.72 / 721.33
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{t^{2}}\mathbf{1}_{t>t_{\min}}$ $\mathcal{C}_{\text{NN}}$ 14.41 17.99 22.74 30.27 32.29 101.12 / 1444.69
$\mathcal{L}_{\text{pow 1}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\text{NN}},$ 14.42 18.20 23.44 32.80 37.17 12.13 / 459.96
$\mathcal{L}_{\text{pow 3}}$ $w_{t}=\frac{1}{(1-t)^{3}}$ $\mathcal{C}_{\text{NN}}$ 14.33 17.96 23.14 32.72 37.23 45.15 / 1189.55
$\mathcal{L}_{\text{mid}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 13.91 18.05 23.51 32.85 37.16 22.34 / 527.53
$\mathcal{L}_{\text{mid-0.2}}$ $w_{t}=\frac{1}{(0.2-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 14.34 18.20 23.41 32.66 36.90 22.67 / 611.83
$\mathcal{L}_{\text{den}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\mathrm{I+NN}},$ 14.42 18.18 23.35 32.58 36.79 19.29 / 717.76
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\mathrm{I+NN}},$ 14.42 17.99 22.78 30.53 34.09 148.95 / 1681.69
$\mathcal{L}_{\text{pow 1}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\mathrm{I+NN}},$ 14.42 18.20 23.44 32.80 37.17 12.64 / 481.42
$\mathcal{L}_{\text{pow 3}}$ $w_{t}=\frac{1}{(1-t)^{3}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.39 18.13 23.36 32.90 37.42 15.32 / 564.80
$\mathcal{L}_{\text{mid}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 13.93 18.04 23.51 32.86 37.22 17.34 / 564.80
$\mathcal{L}_{\text{FM}}+\mathcal{R}_{\text{early}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.18 18.19 23.52 32.97 37.43 10.34 / 323.22
$\mathcal{L}_{\text{FM}}+\mathcal{R}_{\text{late}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.41 17.92 23.43 32.97 37.43 22.81 / 565.44
10-denoisers, $\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\mathrm{NN}}$ 14.42 18.16 23.39 33.02 37.49 9.80 / 265.04
10-denoisers $\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.42 18.18 23.40 33.03 37.50 9.44 / 248.00
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.41 18.22 23.52 32.98 37.44 9.64 / 303.03
[…]
Loss Class PSNR (↑) FID / DINO (↓)
@@
$\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\text{NN}}$ 16.01 20.98 26.25 34.50 37.91 21.13 / 511.97
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{t^{2}}\mathbf{1}_{t>t_{\min}}$ $\mathcal{C}_{\text{NN}}$ 15.88 20.19 24.41 29.35 29.81 87.80 / 1706.42
$\mathcal{L}_{\text{pow 1}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\text{NN}}$ 16.04 21.07 26.44 34.99 38.84 8.36 / 271.09
$\mathcal{L}_{\text{pow 3}}$ $w_{t}=\frac{1}{(1-t)^{3}}$ $\mathcal{C}_{\text{NN}}$ 15.51 20.39 25.79 34.75 38.86 34.39 / 751.94
$\mathcal{L}_{\text{mid}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 15.29 20.82 26.54 35.17 38.95 22.25 / 441.85
$\mathcal{L}_{\text{mid-0.2}}$ $w_{t}=\frac{1}{(0.2-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 15.98 21.12 26.49 34.93 38.59 11.10 / 356.50
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 15.93 20.79 26.12 34.83 38.81 15.48 / 444.89
$\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\mathrm{I+NN}}$ 15.98 20.87 26.10 34.40 38.06 19.02 / 508.06
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{t^{2}}\mathbf{1}_{t>t_{\min}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 15.89 20.26 24.70 31.34 34.67 135.64 / 1368.35
$\mathcal{L}_{\text{pow 1}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\mathrm{I+NN}}$ 16.05 21.07 26.46 35.09 39.03 6.93 / 248.00
$\mathcal{L}_{\text{pow 3}}$ $w_{t}=\frac{1}{(1-t)^{3}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 15.98 20.98 26.40 35.28 39.39 7.10 / 209.17
$\mathcal{L}_{\text{mid}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.80 20.79 26.58 35.22 39.08 21.17 / 419.67
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 16.03 21.08 26.52 35.33 39.34 5.33 / 167.68
(b) FID on 10k tests images.
Figure 7: PSNR and FID for the different losses and parametrizations, CelebA-64, 300 epochs. Models that reach the best PSNR (low values of PSNR difference with respect to standard FM) also reach the lowest FID.
Appendix D Sensitivity of generation to last timesteps
@@
Figure 8: FID comparison when replacing late-time denoising with a pre-trained GS-DRUNet. While the external denoiser slightly improves performance, a gap with the FM baseline remains, suggesting late-time denoising is not the main cause of the discrepancy.
@@
Given the sensitivity of FID to noise, one might wonder whether the poor performance of some denoisers in our toolkit is due to their limited denoising ability at late times (i.e., low noise levels). The following experiment LABEL:{fig:fid_drunet} shows that this is not the case. We generate samples as usual with a Dopri5 scheme, but stop the integration at $t=0.95$ and complete the process with a generic GS-DRUNet pre-trained denoiser from the deepinv library [Tachella et al., 2025], trained on a different dataset. We then compute the FID on 10k generated images and compare it with the original FID. Although plugging in this external denoiser slightly improves the FID, a gap with the FM baseline remains, indicating that late-time denoising is not the only cause of the discrepancy.
Appendix E Inpainting visual results
@@
We display in Figure 9 the reconstructions obtained with our denoisers on a single inpainting task on CelebA-64 with a mask of size $17\times 17$. We set the parameter $\alpha$ in PnP-Flow to $\alpha=0.3$ and the number of iterations to $100$, following the recommendations of the original paper [Martin et al., 2025] and using their public implementation. One can see that only a few methods accurately recover the headband: namely the flow-matching baseline $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$ and the variants $\mathcal{L}_{\text{mid}},\mathcal{C}_{\text{NN}}$ and $\mathcal{L}_{\text{mid}},\mathcal{C}_{\mathrm{I+NN}}$.
@@
(d) $\mathcal{L}_{\text{pow\_1}},\mathcal{C}_{\text{NN}}$
(e) $\mathcal{L}_{\text{pow\_3}},\mathcal{C}_{\text{NN}}$
(f) $\mathcal{L}_{\text{mid}},\mathcal{C}_{\text{NN}}$
(g) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\text{NN}}$
(h) $\mathcal{L}_{\text{den}},\mathcal{C}_{\mathrm{I+NN}}$
(i) $\mathcal{L}_{\text{classic}},\mathcal{C}_{\mathrm{I+NN}}$
(j) $\mathcal{L}_{\text{pow\_1}},\mathcal{C}_{\mathrm{I+NN}}$
(k) $\mathcal{L}_{\text{pow\_3}},\mathcal{C}_{\mathrm{I+NN}}$
(l) $\mathcal{L}_{\text{mid}},\mathcal{C}_{\mathrm{I+NN}}$
(m) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$
@@
Appendix F Perturbation samples

### v2

@@ Appendix B Details on the losses
where $\sigma_{\max}$ is typically around $0.5$. From Remark 1, classical and generative denoisers are equivalent up to the reparameterization $t(\sigma)=\frac{1}{1+\sigma}$. A change of variables in (13) gives:
@@
Appendix C Details on training generation and metrics
All networks are trained with the same random initialization to ensure comparability. For CIFAR-10 we train for 1000 epochs with batch size $128$, and for CelebA-64 and CelebA-128, we train for 300 epochs with batch size $128$. We apply exponential moving average to stabilize training. For CelebA-64 and CelebA-128, we use a U-Net architecture [ronneberger2015unet] as in Ho2020. For CIFAR-10, we adopt the architecture from the torchcfm library [tong2023improving].
The 10 denoisers of Figure 2 are trained independently with 100 epochs each. All samples are generated by solving the ODE using the dopri5 scheme from $t=0$ to $t=1$ with $100$ timesteps; generating samples using a denoiser $D$ means that using the velocity $v(x,t)=\tfrac{D(x,t)-x}{1-t}$.
To measure generation quality, we use the standard Fréchet Inception Distance [heusel2017] with the Inception-V3 features [szegedy2016rethinking]; as recommended by stein2023exposing we also compute it with the DINOv2 embedding.
Appendix D Additional results on CIFAR10 and CelebA-64
Table 2: PSNR and FID for the different losses, to be compared with the standard FM (bottom line) . PSNR computed on 1000 images; FID on 50k train images; CIFAR-10, 1000 epochs.
Loss Class PSNR (↑) FID / DINO (train 50k) (↓)
@@ Appendix D Additional results on CIFAR10 and CelebA-64
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 14.39 18.20 23.50 32.96 37.41 4.89 / 293.84
$\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\text{NN}}$ 14.39 18.20 23.40 32.64 36.85 16.02 / 539.77
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{t^{2}}\mathbf{1}_{t>t_{\min}}$ $\mathcal{C}_{\text{NN}}$ 14.39 18.03 22.89 30.82 33.18 71.07 / 1378.22
$\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.39 18.20 23.39 32.67 36.91 10.29 / 527.02
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{t^{2}}\mathbf{1}_{t>t_{\min}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.39 18.03 22.90 31.00 34.46 116.18 / 1500.15
$\mathcal{L}_{\text{SNR}}$ $w_{t}=\frac{t^{2}}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.37 18.19 23.54 33.05 37.53 3.71 / 181.59
$\mathcal{L}_{\text{high-SNR}}$ $w_{t}=\frac{t^{2}}{(1-t)^{4}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 13.99 17.84 23.24 32.98 37.53 36.66 / 955.25
$\mathcal{L}_{\text{mid 0.2}}$ $w_{t}=\frac{1}{(0.2-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.36 18.22 23.47 32.75 37.06 8.07 / 432.52
$\mathcal{L}_{\text{mid 0.3}}$ $w_{t}=\frac{1}{(0.3-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.26 18.22 23.51 32.81 37.13 8.17 / 361.39
$\mathcal{L}_{\text{mid 0.4}}$ $w_{t}=\frac{1}{(0.4-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.14 18.17 23.54 32.85 37.20 9.80 / 351.951
$\mathcal{L}_{\text{mid 0.5}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 13.93 18.05 23.56 32.91 37.27 9.05 / 313.65
$\mathcal{L}_{\text{mid 0.6}}$ $w_{t}=\frac{1}{(0.6-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.00 17.93 23.56 32.96 37.33 11.25 / 339.08
$\mathcal{L}_{\text{mid 0.7}}$ $w_{t}=\frac{1}{(0.7-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 13.96 17.81 23.51 33.01 37.40 19.39 / 457.23
$\mathcal{L}_{\text{mid 0.8}}$ $w_{t}=\frac{1}{(0.8-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.01 17.74 23.36 33.04 37.46 27.78 / 632.75
$\mathcal{L}_{\text{mid 0.9}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.27 17.80 23.09 33.00 37.50 35.94 / 986.16
$\mathcal{L}_{\text{mid 0.95}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.33 17.93 23.08 32.88 37.48 37.82 / 1097.20
$\mathcal{L}_{\text{FM}}+\mathcal{R}_{\text{[0.1,0.3]}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.32 18.19 23.54 33.01 37.47 6.47 / 279.18
$\mathcal{L}_{\text{FM}}+\mathcal{R}_{\text{[0.3,0.6]}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.36 17.96 23.41 33.01 37.46 24.18 / 601.15
$\mathcal{L}_{\text{FM}}+\mathcal{R}_{\text{[0.6,0.9]}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.38 18.20 23.31 32.86 37.41 29.58 / 564.66
$\mathcal{L}_{\text{FM}}+\mathcal{R}_{\text{[0.4,0.8]}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.38 18.18 23.47 33.01 37.47 16.04 / 459.39
10-denoisers $\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.38 18.15 23.43 32.98 37.45 20.81 / 485.50
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.39 18.21 23.54 33.02 37.47 3.90 / 210.54
Table 3: PSNR and FID for the different losses, to be compared with the standard FM (bottom line). PSNR computed on 1000 test images; FID on 50k train images; CelebA-64, 300 epochs.
Loss Class PSNR (↑) FID / DINO (train 50k) (↓)
@@
$\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\text{NN}}$ 16.01 20.98 26.25 34.50 37.91 19.87 / 518.03
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{t^{2}}\mathbf{1}_{t>t_{\min}}$ $\mathcal{C}_{\text{NN}}$ 15.88 20.19 24.41 29.35 29.81 85.44 / 1708.95
$\mathcal{L}_{\text{mid}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 15.29 20.82 26.54 35.17 38.95 20.06 / 434.22
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 15.93 20.79 26.12 34.83 38.81 14.54 / 448.83
$\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\mathrm{I+NN}}$ 15.98 20.87 26.10 34.40 38.06 18.83 / 513.82
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{t^{2}}\mathbf{1}_{t>t_{\min}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 15.89 20.26 24.70 31.34 34.67 133.93 / 1374.11
$\mathcal{L}_{\text{mid}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.80 20.79 26.58 35.22 39.08 19.37 / 412.68
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 16.03 21.08 26.52 35.33 39.34 4.45 / 164.77
Appendix E Sensitivity of generation to last timesteps
@@
Figure 9: FID (10k-test) comparison on Celeba64 when replacing late-time denoising with a pre-trained GS-DRUNet. While the external denoiser slightly improves performance, a gap with the FM baseline remains, suggesting late-time denoising is not the main cause of the discrepancy.
@@
Given the sensitivity of FID to noise, one might wonder whether the poor performance of some denoisers in our toolkit is due to their limited denoising ability at late times (i.e., low noise levels). The following experiment LABEL:{fig:fid_drunet} shows that this is not the case. We generate samples as usual with a Dopri5 scheme, but stop the integration at $t=0.95$ and complete the process with a generic GS-DRUNet pre-trained denoiser from the deepinv library [tachella2025deepinverse], trained on a different dataset. We then compute the FID on 10k generated images and compare it with the original FID. Although plugging in this external denoiser slightly improves the FID, a gap with the FM baseline remains, indicating that late-time denoising is not the only cause of the discrepancy.
Appendix F Inpainting visual results
@@
We display in Figure 10 the reconstructions obtained with our denoisers on a single inpainting task on CelebA-64 with a mask of size $17\times 17$. We set the parameter $\alpha$ in PnP-Flow to $\alpha=0.3$ and the number of iterations to $100$, following the recommendations of the original paper [martin2024pnp] and using their public implementation. One can see that only a few methods accurately recover the headband: namely the flow-matching baseline $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$ and the variants $\mathcal{L}_{\text{mid}},\mathcal{C}_{\text{NN}}$ and $\mathcal{L}_{\text{mid}},\mathcal{C}_{\mathrm{I+NN}}$.
@@
(d) $\mathcal{L}_{\text{mid}},\mathcal{C}_{\text{NN}}$
(e) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\text{NN}}$
(f) $\mathcal{L}_{\text{den}},\mathcal{C}_{\mathrm{I+NN}}$
(g) $\mathcal{L}_{\text{classic}},\mathcal{C}_{\mathrm{I+NN}}$
(h) $\mathcal{L}_{\text{mid}},\mathcal{C}_{\mathrm{I+NN}}$
(i) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$
@@
Appendix G Ablation on the perturbation experiments Section 5
Despite the checkboard perturbations, we also consider below a range of perturbations defined as follows:
  • Low-frequency - $r$ denotes the pertubration $\delta:=h\star g_{0}$ where $g_{0}$ is a fixed random Gaussian noise and $h$ is a low-pass filter with cutoff frequency $r$;
  • Low-frequency-luminance - $r$ same as above, except $g_{0}$ is shared across the 3 RGB channels (i.e., grayscale noise), and $h$ is a low-pass filter with cutoff frequency $r$.
The higher $r$ is, the more frequency are present (hence evolves from low-freq at $r=0.05$ to high-freq at $r=0.5$) Results on CelebA-128 are in Figure 11, samples are displayed Figures 14 and 15. Results on CIFAR10 are in Figures 16 and 17, samples are displayed Figure 18.
Figure 11: Influence of different perturbations at different generation phases on the FID (10K, test) on CelebA 128.